OpenStudy (anonymous):

The work of a student to solve a set of equations is shown below: Equation 1: x + 2y = 10 Equation 2: 2x + 5y = 15 Step 1: −2(x + 2y) = −2(10) 2x + 5y = 15 Step 2: −2x − 4y = −10 2x + 5y = 15 Step 3: 1y = 5 Step 4: x + 2(1) = 10 Step 5: x = 8 In which step did the student first make an error? Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

−2(10) != −10

OpenStudy (anonymous):

step 1 ?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Yeah a few more if your willing lol

OpenStudy (anonymous):

kk hold on

OpenStudy (anonymous):

Jill wants to solve the following system using the elimination method: y = x + 8 y = 3x − 11 What number should the equation y = x + 8 be multiplied by to eliminate x? −1 −2 −3 3

OpenStudy (anonymous):

I got 3 for this am I right ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Maria wants to solve the following system using the elimination method: 4x + 9y = 59 x + y = 17 What number should the equation x + y = 17 be multiplied by to eliminate y? 4 −4 9 −9

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

if you don't know it thats fine we can just go to the next question

OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: y + z = 4 Equation Q: 3y + 7z = 15 Which of the following steps can be used to eliminate the y term? −3(y + z = 4) 3(y + z = 4) −1(3y + 7z = 15) −3(3y + 7z = 15)

OpenStudy (anonymous):

last one i promise

OpenStudy (anonymous):

I don't want to pressure you, you already helped a lot

OpenStudy (anonymous):

lol thx

OpenStudy (anonymous):

lol its fine. I know its really hard

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

pre algebra

OpenStudy (anonymous):

thats the american school system for ya lol

OpenStudy (anonymous):

orange. you ?

OpenStudy (anonymous):

I used to live there

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

its that bad ?

OpenStudy (anonymous):

I hope it gets easier for you.

OpenStudy (anonymous):

lol me too

OpenStudy (anonymous):

I agree. We should learn things that we actually have to deal with in the real world

OpenStudy (anonymous):

right haha

OpenStudy (anonymous):

im with you on that lol

OpenStudy (anonymous):

nah my parents don't allow me to have social media

OpenStudy (anonymous):

it sucks but I have a kik

OpenStudy (anonymous):

thanks you too

OpenStudy (anonymous):

depends on the person but yeah I guess

OpenStudy (anonymous):

lol i wont

OpenStudy (anonymous):

well you can always talk to me if you like. I used to be shy but when I got out of my shell more things got a whole lot easier

OpenStudy (anonymous):

hold on I forgot the name let me check

OpenStudy (anonymous):

its milalovesu0